OpenStudy (openstudyboy88):

1. A humpback whale can eat 4.4 x 103 plankton a day during feeding season. If feeding season lasts 120 days, find the total number of plankton a humpback would eat by the end of feeding season. Write your answer in scientific notation. 2. The population of Brazil is about 190,000,000, and the population of Sweden is about 9,500,000. How many times greater is the population of Brazil than the population of Sweden? Express your answer in scientific notation.

OpenStudy (sooobored):

\[4.4\times 10^3 \times 120 =?\]

OpenStudy (openstudyboy88):

Idk

OpenStudy (bonnieisflash1.0):